# ContextBench: A Benchmark for Context Retrieval in Coding Agents

### v1

@@ Abstract
LLM-based coding agents have shown strong performance on automated issue resolution benchmarks, yet existing evaluations largely focus on final task success, providing limited insight into how agents retrieve and use code context during problem solving. We introduce ContextBench, a process-oriented evaluation of context retrieval in coding agents. ContextBench consists of 1,136 issue-resolution tasks from 66 repositories across eight programming languages, each augmented with human-annotated gold contexts. We further implement an automated evaluation framework that tracks agent trajectories and measures context recall, precision, and efficiency throughout issue resolution. Using ContextBench, we evaluate four frontier LLMs and five coding agents. Our results show that sophisticated agent scaffolding yields only marginal gains in context retrieval (“The Bitter Lesson” of coding agents), LLMs consistently favor recall over precision, and substantial gaps exist between explored and utilized context. ContextBench augments existing end-to-end benchmarks with intermediate gold-context metrics that unbox the issue-resolution process. These contexts offer valuable intermediate signals for guiding LLM reasoning in software tasks. Data and code are available at: https://cioutn.github.io/context-bench/.
00footnotetext: Equal contribution as co-second authors.00footnotetext: Corresponding author. Co-corresponding author.
@@ 1 Introduction
Large Language Model (LLM) coding agents, exemplified by SWE-agent [23], OpenHands [20], and Agentless [21] have recently demonstrated strong proficiency in real-world software engineering tasks, including bug fixing [11, 16], feature implementation [17, 8], and interaction with terminal-based execution environments [14]. These agents are autonomous systems capable of multi-step reasoning and tool use, enabling them to solve complex, repository-scale tasks in real-world development scenarios.
Limitations of Existing Benchmarks. Advancements in this area are typically assessed on benchmarks such as SWE-bench [11] and SWE-bench Pro [9], which curate tasks from real-world GitHub issues and require agents to submit patches validated by test suites. However, these benchmarks prioritize end-to-end task success rates (e.g., Pass@k) and ignore how coding agents arrive at their solutions. In particular, they overlook the evaluation of context retrieved from large codebases, which underpins the agent’s reasoning. Without verifying intermediate contexts, high success rates may result from trial-and-error or overfitting to specific test cases, yielding coding agents that are unreliable in practical scenarios. We therefore argue for a context-based evaluation system that opens the black box of software tasks and addresses the key question: “How do LLM agents retrieve and use critical code context when resolving software engineering tasks?”
ContextBench: Beyond End-to-End Benchmarking. To address this question, we develop a semi-automated pipeline for curating Gold Contexts, defined as standard sets of repository artifacts that expert developers identify as necessary to resolve a given issue. We begin by sampling high-quality tasks from four widely used benchmarks, i.e., SWE-bench Verified [11], Multi-SWE-bench [25], SWE-PolyBench PB500 [17], and SWE-bench Pro [9]. After deduplication via rule-based matching and embedding-based similarity detection, we select challenging tasks based on three difficulty metrics, i.e., agent solvability; edit scope; and edit dispersion based on gold patches. For each selected task, human experts iteratively annotate and refine the required context in a human-in-the-loop process, ensuring compact and sufficient gold contexts for issue resolution.
@@
Take-Aways. We benchmark four state-of-the-art LLMs, including GPT-5 [19], Claude Sonnet 4.5 [1], Gemini 2.5 Pro [7], and Devstral 2 [15], as well as five widely used coding agents, including mini-SWE-agent, SWE-agent [23], OpenHands [20], Agentless [21], and Prometheus [6]. Our analysis reveals several key findings: ❶ Sophisticated scaffolding does not necessarily lead to better context retrieval performance. Across five coding agents, more complex retrieval scaffolds do not consistently outperform a simple baseline, as shown in 1(a), suggesting potential over-engineering and echoing the Bitter Lesson in coding agents. ❷ ContextBench is challenging for state-of-the-art LLMs. State-of-the-art LLMs struggle to retrieve effective code contexts, often covering relevant information while introducing substantial noise that undermines retrieval precision, as shown in 1(b). ❸ Recall is favored over precision in LLM context retrieval. All evaluated LLMs retrieve broad context to maximize coverage, introducing substantial noise and yielding limited precision and F1 gains. ❹ Balanced retrieval improves performance and cost efficiency. Models that balance retrieval frequency and context granularity achieve higher Pass@1 at lower cost, while aggressive retrieval mainly increases token consumption. ❺ Significant gaps exist between retrieved and utilized context. Agents often inspect gold-relevant code but fail to retain or use it in final patch generation, highlighting consolidation as a key bottleneck.
@@ 2.1 Step 1: Task Deduplication
We curate ContextBench by systematically sampling tasks from four widely used issue resolution benchmarks, i.e., SWE-bench Verified [11] (500 tasks), Multi-SWE-bench [25] (1632 tasks), SWE-PolyBench PB500 [17] (500 tasks), and SWE-bench Pro [9] (1865 tasks), totaling 4,497 tasks. These benchmarks offer high-quality task sources curated from real-world GitHub repositories and issues, with strong reproducibility guarantees and multilingual coverage across Python, Java, JavaScript, TypeScript, Go, Rust, C, and C++.
@@ 2.3 Step 3: Expert Annotation
11footnotetext: https://github.com/langchain-ai/langchain/pull/4009
@@ 3 Can Coding Agents Retrieve Effective Contexts to Resolve Issues?
3.1 RQ1: Benchmarking Coding Agents
Table 2 benchmarks five coding agents that are open-sourced, actively maintained, and representative of major paradigms in context retrieval. Specifically, we assess a standard baseline that retrieves context through simple bash commands (i.e., min-SWE-agent [23]) and four state-of-the-art scaffolds, including embedding-based semantic retrieval (i.e., Agentless [21]), file-system navigation via specialized interfaces (i.e., SWE-agent [23], OpenHands [20]), and graph-based repository retrieval (i.e., Prometheus [6]). All agents are evaluated with an advanced LLM backbone (i.e., GPT-5 [19]). Detailed experiment settings are listed in Section C.2.
3.2 RQ2: Benchmarking Large Language Models
Table 3 benchmarks four state-of-the-art LLMs, including GPT-5 [19], Claude Sonnet 4.5 [1], Gemini 2.5 Pro [7], and Devstral 2 [15], in retrieving relevant code contexts from large-scale codebases when resolving issues using the standard coding agent scaffold (i.e., mini-SWE-agent [23]). Detailed experiment settings are listed in Section C.2.
@@ 3.3 RQ3: Analysis of Context Retrieval Patterns
3.4 RQ4: Analysis of Context Retrieval Dynamics
@@ 4 Related Work
Repository-Level Code Evaluation Benchmarks. Repository-level benchmarks ground evaluation in real-world codebases and executable test suites, making them the de facto standard for assessing end-to-end issue resolution. SWE-bench and its variants (Lite, Verified, multilingual extensions) evaluate whether a system can produce a test-passing patch under realistic repository settings [23, 11], while complementary benchmarks such as RepoBench [13] and SWE-PolyBench [17] emphasize cross-file completion or structure-aware analysis. Despite their realism, these benchmarks are predominantly outcome-driven: they largely measure final task success, offering limited visibility into whether agents succeed by identifying the right information or by compensating through extensive exploration [6]. Our work complements outcome-only evaluation by (i) introducing verified gold contexts and (ii) enabling dynamic, process-level evaluation of context retrieval beyond final outcomes.
Coding Agents. LLM-based coding agents address repository-level issue resolution through either pipeline-based or agentic designs. Pipeline-based systems decompose tasks into fixed stages with controlled context selection [21, 26, 18, 10, 24], agentic systems adopt open-ended interaction loops in which agents dynamically search, inspect, and modify codebases via tool calls [23, 20, 22, 4, 6]. Despite architectural diversity, context retrieval remains a shared bottleneck. Existing benchmarks primarily evaluate end-to-end success, making it difficult to compare how agents retrieve and manage code context. ContextBench fills this gap by providing a retrieval-centric benchmark that directly evaluates context acquisition behavior, independent of final patch correctness.
Trajectory Analysis and Failure Diagnosis. Recent work analyzes agent trajectories to diagnose failure modes via manual inspection, error taxonomies, or large-scale statistics over execution logs [27, 12, 5, 2, 3]. Such analyses have improved understanding of where agents fail (e.g., reasoning deadlocks, tool misuse, runtime exceptions), and some studies report that context-related errors are frequent in repository-level settings. However, trajectory analysis is often labor-intensive, difficult to standardize across frameworks, and typically provides post hoc explanations rather than instance-level supervision. ContextBench addresses this gap by offering a scalable, standardized evaluation layer: verified gold contexts and automatic metrics that quantify retrieval effectiveness without requiring manual reading of long trajectories, while still enabling connections between intermediate retrieval behavior and end-to-end outcomes.
@@ C.1 Model Variations within mini-SWE-Agent
During the evaluation, we observed that DevStral2 occasionally exhibited unexpected behaviors that deviated from the standardized output protocol. These anomalies are discussed in detail in Appendix L, where we provide the corresponding raw interaction traces and highlight the specific issues encountered.
@@
Strict Output Format Constraint **Requirements:**
List ONLY the specific code sections from source files that you examined to create your patch.
**Format (strictly follow):**
Each entry must include:
@@
**Example:**
File: /testbed/src/core/handler.ext
@@
**DO NOT include:**
- Explanations, reasoning, or commentary
@@ C.2 Cross-Agent Comparison for the Same Model
Pre-Submission Context Enforcement Specification in SWE-Agent Required Format: In the next response, the agent must: 1. Output a <PATCH_CONTEXT> block listing all source files and line ranges examined. 2. Invoke the submit function. Format for <PATCH_CONTEXT>:
@@
Rules: List only source files that were viewed or edited (exclude test files). Include the exact line ranges examined. Do not include code snippets, explanations, or commentary. This step is mandatory and cannot be skipped.
@@
Context Extraction Protocol in Openhands (system_prompt_context_aware.j2) Explore-context marking protocol. When exploring source code context (e.g., reading files or printing line ranges), the agent is required to emit a machine-parseable explore-context block before the corresponding command. <EXPLORE_CONTEXT> File: /absolute/path/to/file.ext Lines: <start>-<end> </EXPLORE_CONTEXT> The explore-context block must contain one or more file-range entries with absolute paths and positive line indices. It is only included when the command prints source code content to stdout and must not be used for metadata-only commands (e.g., ls, grep, git status). Patch Context Submission Requirement Before finalizing the submission, the agent must provide the code context used for patch generation. Format (strictly enforced): THOUGHT: Providing code context used for patch generation. <PATCH_CONTEXT> File: <absolute_file_path> Lines: <start_line>-<end_line> File: <absolute_file_path> Lines: <start_line>-<end_line> </PATCH_CONTEXT> Only absolute file paths and line ranges are permitted inside the <PATCH_CONTEXT> block. Explanations, comments, debugging output, test code, or any non-code content are strictly disallowed. Unlike mini-SWE-agent and SWE-Agent, Openhands employs a different prompt format. Nevertheless, this variation in prompt design does not preclude the implementation of an equivalent context extraction mechanism. By adapting the extraction protocol to accommodate the agent-specific prompt structure, we ensure that the resulting <PATCH_CONTEXT> outputs remain consistent, structured, and machine-verifiable, thereby preserving the integrity and comparability of the context extraction process across heterogeneous agents.
@@ I.3 Case 3: OpenHands Agent – Cross-Context Exploration Gap
Appendix J Failure Taxonomy for Agents
Below, we present the clustering and analysis of error labels for each agent.
Figure 6: : Taxonomy of failure types
Figure 7: : Failure taxonomy for Agentless
Figure 8: : Failure taxonomy for Openhands
Figure 9: : Failure taxonomy for Prometheus
Appendix K Cross-Language Radar Distributions for File/Block/Line Metrics
@@
Appendix L Potential Data Contamination and Protocol Misalignment in Devstral2

### v2

@@ Abstract
LLM-based coding agents have shown strong performance on automated issue resolution benchmarks, yet existing evaluations largely focus on final task success, providing limited insight into how agents retrieve and use code context during problem solving. We introduce ContextBench, a process-oriented evaluation of context retrieval in coding agents. ContextBench consists of 1,136 issue-resolution tasks from 66 repositories across eight programming languages, each augmented with human-annotated gold contexts. We further implement an automated evaluation framework that tracks agent trajectories and measures context recall, precision, and efficiency throughout issue resolution. Using ContextBench, we evaluate four frontier LLMs and five coding agents. Our results show that sophisticated agent scaffolding yields only marginal gains in context retrieval (“The Bitter Lesson” of coding agents), LLMs consistently favor recall over precision, and substantial gaps exist between explored and utilized context. ContextBench augments existing end-to-end benchmarks with intermediate gold-context metrics that unbox the issue-resolution process. These contexts offer valuable intermediate signals for guiding LLM reasoning in software tasks.
@@ 1 Introduction
Large Language Model (LLM) coding agents, exemplified by SWE-agent Yang et al. (2024), OpenHands Wang et al. (2025), and Agentless Xia et al. (2024) have recently demonstrated strong proficiency in real-world software engineering tasks, including bug fixing Jimenez et al. (2024); Pan et al. (2025), feature implementation Rashid et al. (2025); Deng et al. (2025a), and interaction with terminal-based execution environments Merrill et al. (2026). These agents are autonomous systems capable of multi-step reasoning and tool use, enabling them to solve complex, repository-scale tasks in real-world development scenarios.
Limitations of Existing Benchmarks. Advancements in this area are typically assessed on benchmarks such as SWE-bench Jimenez et al. (2024) and SWE-bench Pro Deng et al. (2025b), which curate tasks from real-world GitHub issues and require agents to submit patches validated by test suites. However, these benchmarks prioritize end-to-end task success rates (e.g., Pass@k) and ignore how coding agents arrive at their solutions. In particular, they overlook the evaluation of context retrieved from large codebases, which underpins the agent’s reasoning. Without verifying intermediate contexts, high success rates may result from trial-and-error or overfitting to specific test cases, yielding coding agents that are unreliable in practical scenarios. We therefore argue for a context-based evaluation system that opens the black box of software tasks and addresses the key question: “How do LLM agents retrieve and use critical code context when resolving software engineering tasks?”
ContextBench: Beyond End-to-End Benchmarking. To address this question, we develop a semi-automated pipeline for curating Gold Contexts, defined as standard sets of repository artifacts that expert developers identify as necessary to resolve a given issue. We begin by sampling high-quality tasks from four widely used benchmarks, i.e., SWE-bench Verified Jimenez et al. (2024), Multi-SWE-bench Zan et al. (2025a), SWE-PolyBench PB500 Rashid et al. (2025), and SWE-bench Pro Deng et al. (2025b). After deduplication via rule-based matching and embedding-based similarity detection, we select challenging tasks based on three difficulty metrics, i.e., agent solvability; edit scope; and edit dispersion based on gold patches. For each selected task, human experts iteratively annotate and refine the required context in a human-in-the-loop process, ensuring compact and sufficient gold contexts for issue resolution.
@@
Take-Aways. We benchmark four state-of-the-art LLMs, including GPT-5 Singh et al. (2025), Claude Sonnet 4.5 Anthropic (2025), Gemini 2.5 Pro Comanici et al. (2025), and Devstral 2 Mistral AI (2025), as well as five widely used coding agents, including mini-SWE-agent, SWE-agent Yang et al. (2024), OpenHands Wang et al. (2025), Agentless Xia et al. (2024), and Prometheus Chen et al. (2025c). Our analysis reveals several key findings: ❶ Sophisticated scaffolding does not necessarily lead to better context retrieval performance. Across five coding agents, more complex retrieval scaffolds do not consistently outperform a simple baseline, as shown in 1(a), suggesting potential over-engineering and echoing “The Bitter Lesson” of AI research. ❷ ContextBench is challenging for state-of-the-art LLMs. State-of-the-art LLMs struggle to retrieve effective code contexts, often covering relevant information while introducing substantial noise that undermines retrieval precision, as shown in 1(b). ❸ Recall is favored over precision in LLM context retrieval. All evaluated LLMs retrieve broad context to maximize coverage, introducing substantial noise and yielding limited precision and F1 gains. ❹ Balanced retrieval improves performance and cost efficiency. Models that balance retrieval frequency and context granularity achieve higher Pass@1 at lower cost, while aggressive retrieval mainly increases token consumption. ❺ Significant gaps exist between retrieved and utilized context. Agents often inspect gold-relevant code but fail to retain or use it in final patch generation, highlighting consolidation as a key bottleneck.
@@ 2.1 Step 1: Task Deduplication
We curate ContextBench by systematically sampling tasks from four widely used issue resolution benchmarks, i.e., SWE-bench Verified Jimenez et al. (2024) (500 tasks), Multi-SWE-bench Zan et al. (2025a) (1632 tasks), SWE-PolyBench PB500 Rashid et al. (2025) (500 tasks), and SWE-bench Pro Deng et al. (2025b) (1865 tasks), totaling 4,497 tasks. These benchmarks offer high-quality task sources curated from real-world GitHub repositories and issues, with strong reproducibility guarantees and multilingual coverage across Python, Java, JavaScript, TypeScript, Go, Rust, C, and C++.
@@ 3 Can Coding Agents Retrieve Effective Contexts to Resolve Issues?
3.1 RQ1: Benchmarking Coding Agents
Table 2 benchmarks five coding agents that are open-sourced, actively maintained, and representative of major paradigms in context retrieval. Specifically, we assess a standard baseline that retrieves context through simple bash commands (i.e., min-SWE-agent Yang et al. (2024)) and four state-of-the-art scaffolds, including embedding-based semantic retrieval (i.e., Agentless Xia et al. (2024)), file-system navigation via specialized interfaces (i.e., SWE-agent Yang et al. (2024), OpenHands Wang et al. (2025)), and graph-based repository retrieval (i.e., Prometheus Chen et al. (2025c)). All agents are evaluated with an advanced LLM backbone (i.e., GPT-5 Singh et al. (2025)). Detailed experiment settings are listed in Section C.2.
3.2 RQ2: Benchmarking Large Language Models
Table 3 benchmarks four state-of-the-art LLMs, including GPT-5 Singh et al. (2025), Claude Sonnet 4.5 Anthropic (2025), Gemini 2.5 Pro Comanici et al. (2025), and Devstral 2 Mistral AI (2025), in retrieving relevant code contexts from large-scale codebases when resolving issues using the standard coding agent scaffold (i.e., mini-SWE-agent Yang et al. (2024)). Detailed experiment settings are listed in Section C.2.
@@ 3.4 RQ4: Analysis of Context Retrieval Dynamics
3.5 RQ5: Analysis of Gold Context Robustness
@@ 4 Related Work
Repository-Level Code Evaluation Benchmarks. Repository-level benchmarks ground evaluation in real-world codebases and executable test suites, making them the de facto standard for assessing end-to-end issue resolution. SWE-bench and its variants (Lite, Verified, multilingual extensions Zan et al. (2025b), long-horizon tasks Deng et al. (2025c), live environments Zhang et al. (2025), goal-oriented settings Yang et al. (2025a), and freelance scenarios Miserendino et al. (2025)) evaluate whether a system can produce a test-passing patch under realistic repository settings Yang et al. (2024); Jimenez et al. (2024), while complementary benchmarks such as RepoBench Liu et al. (2023) and SWE-PolyBench Rashid et al. (2025) emphasize cross-file completion or structure-aware analysis. Despite their realism, these benchmarks are predominantly outcome-driven: they largely measure final task success, offering limited visibility into whether agents succeed by identifying the right information or by compensating through extensive exploration Chen et al. (2025c). Our work complements outcome-only evaluation by (i) introducing verified gold contexts and (ii) enabling dynamic, process-level evaluation of context retrieval beyond final outcomes.
Coding Agents. LLM-based coding agents address repository-level issue resolution through either pipeline-based or agentic designs. Pipeline-based systems decompose tasks into fixed stages with controlled context selection Xia et al. (2024); Zhang et al. (2024); Ruan et al. (2024); Gao et al. (2025); Yang et al. (2025b), agentic systems adopt open-ended interaction loops in which agents dynamically search, inspect, and modify codebases via tool calls Yang et al. (2024); Wang et al. (2025); Xia et al. (2025); Chen et al. (2025a, c). Despite architectural diversity, context retrieval remains a shared bottleneck. Existing benchmarks primarily evaluate end-to-end success, making it difficult to compare how agents retrieve and manage code context. ContextBench fills this gap by providing a retrieval-centric benchmark that directly evaluates context acquisition behavior, independent of final patch correctness.
Trajectory Analysis and Failure Diagnosis. Recent work analyzes agent trajectories to diagnose failure modes via manual inspection, error taxonomies, or large-scale statistics over execution logs Zhu et al. (2025); Liu et al. (2025); Chen et al. (2025b); Bouzenia and Pradel (2025); Cemri et al. (2025); Pathak et al. (2025). Such analyses have improved understanding of where agents fail (e.g., reasoning deadlocks, tool misuse, runtime exceptions), and some studies report that context-related errors are frequent in repository-level settings. However, trajectory analysis is often labor-intensive, difficult to standardize across frameworks, and typically provides post hoc explanations rather than instance-level supervision. ContextBench addresses this gap by offering a scalable, standardized evaluation layer: verified gold contexts and automatic metrics that quantify retrieval effectiveness without requiring manual reading of long trajectories, while still enabling connections between intermediate retrieval behavior and end-to-end outcomes.
@@ C.1.4 Cross-Model Evaluation Results
During the evaluation, we observed that DevStral2 occasionally exhibited unexpected behaviors that deviated from the standardized output protocol. These anomalies are discussed in detail in Appendix K, where we provide the corresponding raw interaction traces and highlight the specific issues encountered.
@@
Requirements List ONLY the specific code sections from source files that you examined to create your patch. Format (strictly follow) Each entry must include:
@@
Example File: /testbed/src/core/handler.ext
@@
DO NOT Include - Explanations, reasoning, or commentary
@@ C.2.1 Context Extraction Protocol Extension
C.2.2 SWE-Agent Implementation
Pre-Submission Context Enforcement Specification in SWE-Agent Required Format: In the next response, the agent must: 1. Output a <PATCH_CONTEXT> block listing all source files and line ranges examined. 2. Invoke the submit function. Format for <PATCH_CONTEXT>
@@ C.2.2 SWE-Agent Implementation
Rules List only source files that were viewed or edited (exclude test files). Include the exact line ranges examined. Do not include code snippets, explanations, or commentary. This step is mandatory and cannot be skipped.
@@
C.2.3 Openhands Implementation
Context Extraction Protocol in Openhands (system_prompt_context_aware.j2) Explore-context marking protocol. When exploring source code context (e.g., reading files or printing line ranges), the agent is required to emit a machine-parseable explore-context block before the corresponding command. Explore Context Format <EXPLORE_CONTEXT> File: /absolute/path/to/file.ext Lines: <start>-<end> </EXPLORE_CONTEXT> The explore-context block must contain one or more file-range entries with absolute paths and positive line indices. It is only included when the command prints source code content to stdout and must not be used for metadata-only commands (e.g., ls, grep, git status). Patch Context Submission Requirement Before finalizing the submission, the agent must provide the code context used for patch generation. Format (strictly enforced): THOUGHT: Providing code context used for patch generation. <PATCH_CONTEXT> File: <absolute_file_path> Lines: <start_line>-<end_line> File: <absolute_file_path> Lines: <start_line>-<end_line> </PATCH_CONTEXT> Only absolute file paths and line ranges are permitted inside the <PATCH_CONTEXT> block. Explanations, comments, debugging output, test code, or any non-code content are strictly disallowed. Unlike mini-SWE-agent and SWE-Agent, Openhands employs a different prompt format. Nevertheless, this variation in prompt design does not preclude the implementation of an equivalent context extraction mechanism. By adapting the extraction protocol to accommodate the agent-specific prompt structure, we ensure that the resulting <PATCH_CONTEXT> outputs remain consistent, structured, and machine-verifiable, thereby preserving the integrity and comparability of the context extraction process across heterogeneous agents.
C.2.4 Agentless Enhancements
Bug Fixes and Compatibility.
@@ I.3 Case 3: OpenHands Agent – Cross-Context Exploration Gap
Appendix J Cross-Language Radar Distributions for File/Block/Line Metrics
@@
Appendix K Potential Data Contamination and Protocol Misalignment in Devstral2